What is the percent error in each of the following approximation of ? (a) 22/7 (b)
355/113
1
Expert's answer
2021-09-06T11:35:25-0400

(a) The error is


"\frac{\bigg(\frac{22}{7}-\pi\bigg)}{\pi}·100\%=0.04\%."

(b) The error is


"\frac{\bigg(\frac{355}{113}-\pi\bigg)}{\pi}·100\%=(8.49·10^{-6})\%."